A body is projected vertically up. At certain height h above the ground its PE and KEare in the ratio 1:4. At what height above the ground PEand KE will be in the ratio 4:1?
(a)4h
(b)h/4
(c)5h
(d)h/5

Grade:11

1 Answers

Piyush Kumar Maurya
5 years ago
PE/KE = ¼     - given
4PE = KE
So let at height h'  PE and KE are in the ratio 4:1
So for h' : PE/4 = KE 
 
PE + KE = constant
(PE)+ (KE)= (PE)h' + (KE)h' 
(PE)h + (4PE )= (PE)h' + ( PE/4)h'
(5PE)= (5/4PE)h'
5mgh = 5/4 mgh'
h' = 4h